Kindly provide me just not the answer bt the solution of problem below.
 
One litre flask contains air, water and small amount of liquid water at a pressure of 200 mm Hg. If this is connected to another one litre evacuated flask, what will be the final pressure of the gas mixture at equilibrium. Assume the temperature to be 50 degrees.
(Aqueous tension at 50 degrees=93 mm Hg)

Grade:12th pass

1 Answers

Arun
5 years ago
 
Here the initial pressure is p1=200mm+93mm=293mm
let the final pressure be p2
now... convert the temperature to kelvin scale,t1=t2=323
p1*v1/t1=p2*v2/t2
293*1/323=p2*2/323
solving we get
p2=293/2
p2=146.5mm